# Spiking Neural Networks for Continuous Control via End-to-End Model-Based Learning

### v2

@@ 4 Discussion and Conclusion
Looking ahead, we aim to extend this architecture to reinforcement learning tasks, which introduce challenges in exploration, long-term credit assignment, and reward sparsity. Our Pred-Control SNN is well-suited for this transition due to its modularity and internal state dynamics. However, the reliance on BPTT underscores the importance of exploring alternative learning methods. Three complementary directions appear especially promising. First, online training approaches such as e-prop (Bellec et al.,, 2020) and forward propagation through time (Yin et al.,, 2023) provide more memory-efficient schemes for recurrent spiking networks. Second, noise-driven credit assignment methods such as node and weight perturbation (Züge et al.,, 2023) may offer biologically inspired, event-driven alternatives that scale more gracefully than full backpropagation. Third, benchmarking these methods on high-dimensional continuous control remains a crucial test. MuJoCo-style continuous control environments, already used in recent SNN-based reinforcement learning studies (Tang et al.,, 2021; Zanatta et al.,, 2024), provide a natural next step for assessing whether these algorithms can scale to complex, high-DoF robotic tasks. In parallel, we also intend to explore the generative power of model-based reinforcement learning algorithms adapted to the spiking domain, where predictive components may further reduce reliance on expensive gradient propagation. Such work will be crucial for assessing whether spiking systems can rise to the challenges of real-world robotics, offering not only compact and energy-efficient control but also biologically inspired online adaptability.
@@ Appendix A Learning Rate $\alpha$
In this experiment, we investigate suitable learning rates for two separately trained networks: the policy network $\bm{\pi}$ and the prediction (dynamics) model $\bm{\upsilon}$. Their respective learning rates are denoted as $\alpha_{\bm{\pi}}$ and $\alpha_{\bm{\upsilon}}$. We treat both as independent hyperparameters and sample their values logarithmically between $10^{-5}$ and $10^{-1}$. Training is performed using the Adam optimizer Kingma and Ba, (2014) for both networks, due to its adaptive update rules and strong empirical performance across a wide range of tasks.
@@ Appendix B Surrogate Gradient Function $f^{\prime}(U)$
Training spiking neural networks (SNNs) with backpropagation relies on surrogate gradients $f^{\prime}(U)$ to approximate the non-differentiable spike function during the backward pass (Neftci et al.,, 2019; Zenke and Vogels,, 2021) (see Figure 7, top). Several surrogate functions have been proposed, each differing in shape and defined by key parameters: the steepness $\beta$, which controls the sharpness of the response around threshold, and the scaling factor $\gamma$, which sets the overall gradient amplitude.
@@
  • SuperSpike (Zenke and Ganguli,, 2018), a heavy-tailed approximation with smoother decay
  • GaussianSpike (Yin et al.,, 2021), a mixture-based method with localized support
@@ Appendix C Empirical Initialization
Spiking neural networks are highly sensitive to weight initialization, as only specific parameter regimes yield sufficient activity and stable learning. We adopt a fluctuation-driven initialization scheme (Rossbroich et al.,, 2022), which incorporates the neuron time constants $\tau_{\text{mem}}$ and $\tau_{\text{syn}}$ into the weight scaling rules. This method aims to place the network in a dynamic regime where neurons spike regularly but not excessively, facilitating gradient-based learning. Unlike standard methods such as Kaiming or Xavier initialization—which do not account for temporal filtering—this scheme is better suited to SNNs with explicit dynamics.
@@
Although our networks are relatively shallow (3 layers), they are unrolled over many time steps during task execution, producing significant temporal depth. This contrasts with feedforward SNNs used in classification tasks (Rossbroich et al.,, 2022), where stability was more closely linked to performance. Here, we observe that well-performing configurations are not always those with the lowest loss or the highest activity levels, indicating that different metrics may emphasize different aspects of task behavior.
We fix the target membrane statistics to $\mu_{U}=0$ and $\sigma_{U}=1$ for all initializations. The full derivation of how $\nu$, $\tau_{\text{mem}}$, and $\tau_{\text{syn}}$ influence the initialization scale is provided in (Rossbroich et al.,, 2022).
@@
Interestingly, the configurations with the lowest policy and prediction losses are not those with the highest time-on-target. This mismatch highlights that minimizing error signals does not always correspond to robust goal-directed behavior—reaching the target briefly and maintaining presence there are qualitatively different challenges. Moreover, some parameter settings that would be considered unstable by classical criteria (e.g., low spike coverage or imbalanced layer activity) still yield strong task performance. This stands in partial contrast to prior work on SNN classification (Rossbroich et al.,, 2022), where stable activity and balanced firing were stronger predictors of success. Our findings suggest that the link between stability and performance is more nuanced in closed-loop, temporally extended control tasks.
@@ D.1 Recurrent Connections
A value of $\rho=0.9$ implies a weakly recurrent layer (mostly driven by external input), while $\rho=0.1$ yields a strongly recurrent layer (dominated by recurrent feedback). We sweep $\rho$ from 0.1 to 0.9 in steps of 0.2, using the fluctuation-driven initialization strategy of Rossbroich et al., (2022).
@@ Appendix E Training Procedure
1:Inputs: $\bm{\upsilon}(\hat{\bm{s^{\prime}}}|\bm{s},\bm{u},\bm{\theta_{\upsilon}})$, $\bm{\pi}(\bm{u}|\bm{s},\bm{s}^{*},\bm{\theta_{\pi}})$, environment $(\bm{s^{\prime}},\bm{s}^{*\prime}|\bm{u})$, memory capacity $M$
@@ Appendix F Prediction Network $\bm{\upsilon}$ Training
1:Inputs: $\bm{\upsilon}(\hat{\bm{s}}^{\prime}|\bm{s},\bm{u},\bm{\theta_{\upsilon}})$, memory buffer, $\mathcal{L}_{\bm{\upsilon}}$, $\alpha_{\bm{\upsilon}}$, warmup steps $T_{\text{warm}}$, unroll steps $T_{\text{unroll}}$, teacher forcing prob. $p_{\text{tf}}$
@@ Appendix G Policy Network $\bm{\pi}$ Training
The policy network $\bm{\pi}$ generates control signals that guide the robot arm toward a desired end-effector position. Its architecture mirrors the prediction model, consisting of two spiking populations followed by a leaky integrator layer, but it does not include recurrent connections (see subsection D.1). The network receives the current state $\bm{s}_{t}$ and the target state $\bm{s}_{t}^{*}$ as input and outputs a continuous action $\bm{u}_{t}$.
@@
1:Inputs: $\bm{\pi}(\bm{u}|\bm{s},\bm{s}^{*},\bm{\theta_{\pi}})$, $\bm{\upsilon}(\hat{\bm{s}}^{\prime}|\bm{s},\bm{u},\bm{\theta_{\upsilon}})$, memory buffer, $\mathcal{L}_{\bm{\pi}}$, $\alpha_{\bm{\pi}}$, warmup steps $T_{\text{warm}}$, unroll steps $T_{\text{unroll}}$
@@ Appendix I Adaptive Leaky Integrate-and-Fire Neuron Model
In the following subsections, we examine how each mechanism influences the behavior of single ALIF neurons and their firing responses to tonic and noisy inputs, as well as their impact on learning the 2D control task. To allow both $\vartheta_{a}$ and $\vartheta_{b}$ to contribute meaningfully, in the control experiments we use an unroll window $T_{\text{unroll}}=40$ environment steps (= 240 SNN steps) for both the prediction and the policy model.
@@ I.3 Combined Behavior
Overall, low-rank factorization provides a practical method for scaling SNN models while controlling memory and compute requirements. This approach also aligns with biological findings suggesting that population activity often resides on low-dimensional manifolds (Averbeck et al.,, 2006; Churchland et al.,, 2012).
@@
Appendix O Hyperparameters
@@
Unroll steps (prediction/policy) $T_{\text{unroll}}^{\upsilon},T_{\text{unroll}}^{\pi}$ 10 / 40
@@
Synaptic time constant $\tau_{\text{syn}}$ 2 ms (learned)
Adaptation time constant $\tau_{\text{ada}}$ 0.1 (learned)

### v3

@@ 4 Discussion and Conclusion
Looking ahead, we aim to extend this architecture to reinforcement learning tasks, which introduce challenges in exploration, long-term credit assignment, and reward sparsity. Our Pred-Control SNN is well-suited for this transition due to its modularity and internal state dynamics. However, the reliance on BPTT underscores the importance of exploring alternative learning methods. Three complementary directions appear especially promising. First, online training approaches such as e-prop (Bellec et al., 2020) and forward propagation through time (Yin et al., 2023) provide more memory-efficient schemes for recurrent spiking networks. Second, noise-driven credit assignment methods such as node and weight perturbation (Züge et al., 2023) may offer biologically inspired, event-driven alternatives that scale more gracefully than full backpropagation. Third, benchmarking these methods on high-dimensional continuous control remains a crucial test. MuJoCo-style continuous control environments, already used in recent SNN-based reinforcement learning studies (Tang et al., 2021; Zanatta et al., 2024), provide a natural next step for assessing whether these algorithms can scale to complex, high-DoF robotic tasks. In parallel, we also intend to explore the generative power of model-based reinforcement learning algorithms adapted to the spiking domain, where predictive components may further reduce reliance on expensive gradient propagation. Such work will be crucial for assessing whether spiking systems can rise to the challenges of real-world robotics, offering not only compact and energy-efficient control but also biologically inspired online adaptability.
@@ Appendix A Learning Rate $\alpha$
In this experiment, we investigate suitable learning rates for two separately trained networks: the policy network $\bm{\pi}$ and the prediction (dynamics) model $\bm{\upsilon}$. Their respective learning rates are denoted as $\alpha_{\bm{\pi}}$ and $\alpha_{\bm{\upsilon}}$. We treat both as independent hyperparameters and sample their values logarithmically between $10^{-5}$ and $10^{-1}$. Training is performed using the Adam optimizer Kingma and Ba (2014) for both networks, due to its adaptive update rules and strong empirical performance across a wide range of tasks.
@@ Appendix B Surrogate Gradient Function $f^{\prime}(U)$
Training spiking neural networks (SNNs) with backpropagation relies on surrogate gradients $f^{\prime}(U)$ to approximate the non-differentiable spike function during the backward pass (Neftci et al., 2019; Zenke and Vogels, 2021) (see Figure 7, top). Several surrogate functions have been proposed, each differing in shape and defined by key parameters: the steepness $\beta$, which controls the sharpness of the response around threshold, and the scaling factor $\gamma$, which sets the overall gradient amplitude.
@@
  • SuperSpike (Zenke and Ganguli, 2018), a heavy-tailed approximation with smoother decay
  • GaussianSpike (Yin et al., 2021), a mixture-based method with localized support
@@ Appendix C Empirical Initialization
Spiking neural networks are highly sensitive to weight initialization, as only specific parameter regimes yield sufficient activity and stable learning. We adopt a fluctuation-driven initialization scheme (Rossbroich et al., 2022), which incorporates the neuron time constants $\tau_{\text{mem}}$ and $\tau_{\text{syn}}$ into the weight scaling rules. This method aims to place the network in a dynamic regime where neurons spike regularly but not excessively, facilitating gradient-based learning. Unlike standard methods such as Kaiming or Xavier initialization—which do not account for temporal filtering—this scheme is better suited to SNNs with explicit dynamics.
@@
Although our networks are relatively shallow (3 layers), they are unrolled over many time steps during task execution, producing significant temporal depth. This contrasts with feedforward SNNs used in classification tasks (Rossbroich et al., 2022), where stability was more closely linked to performance. Here, we observe that well-performing configurations are not always those with the lowest loss or the highest activity levels, indicating that different metrics may emphasize different aspects of task behavior.
We fix the target membrane statistics to $\mu_{U}=0$ and $\sigma_{U}=1$ for all initializations. The full derivation of how $\nu$, $\tau_{\text{mem}}$, and $\tau_{\text{syn}}$ influence the initialization scale is provided in (Rossbroich et al., 2022).
@@
Interestingly, the configurations with the lowest policy and prediction losses are not those with the highest time-on-target. This mismatch highlights that minimizing error signals does not always correspond to robust goal-directed behavior—reaching the target briefly and maintaining presence there are qualitatively different challenges. Moreover, some parameter settings that would be considered unstable by classical criteria (e.g., low spike coverage or imbalanced layer activity) still yield strong task performance. This stands in partial contrast to prior work on SNN classification (Rossbroich et al., 2022), where stable activity and balanced firing were stronger predictors of success. Our findings suggest that the link between stability and performance is more nuanced in closed-loop, temporally extended control tasks.
@@ D.1 Recurrent Connections
A value of $\rho=0.9$ implies a weakly recurrent layer (mostly driven by external input), while $\rho=0.1$ yields a strongly recurrent layer (dominated by recurrent feedback). We sweep $\rho$ from 0.1 to 0.9 in steps of 0.2, using the fluctuation-driven initialization strategy of Rossbroich et al. (2022).
@@ Appendix E Training Procedure
1:Inputs: $\bm{\upsilon}(\hat{\bm{s}}_{t+1}|\bm{s}_{t},\bm{u}_{t},\bm{\theta_{\upsilon}})$, $\bm{\pi}(\bm{u}_{t}|\bm{s}_{t},\bm{s}_{t}^{*},\bm{\theta_{\pi}})$, environment $(\bm{s}_{t+1},\bm{s}^{*}_{t+1}|\bm{u}_{t})$, memory capacity $M$
@@ Appendix F Prediction Network $\bm{\upsilon}$ Training
1:Inputs: $\bm{\upsilon}(\hat{\bm{s}}_{t+1}|\bm{s}_{t},\bm{u}_{t},\bm{\theta_{\upsilon}})$, memory buffer, $\mathcal{L}_{\bm{\upsilon}}$, $\alpha_{\bm{\upsilon}}$, warmup steps $T_{\text{warm}}$, unroll steps $T_{\text{unroll}}$, teacher forcing prob. $p_{\text{tf}}$
@@ Appendix G Policy Network $\bm{\pi}$ Training
The policy network $\bm{\pi}$ generates control signals that guide the robot arm toward a desired end-effector position. Its architecture mirrors the prediction model, consisting of two spiking populations followed by a leaky integrator layer, but it does not include recurrent connections (see Section D.1). The network receives the current state $\bm{s}_{t}$ and the target state $\bm{s}_{t}^{*}$ as input and outputs a continuous action $\bm{u}_{t}$.
@@
1:Inputs: $\bm{\pi}(\bm{u}_{t}|\bm{s}_{t},\bm{s}^{*}_{t},\bm{\theta_{\pi}})$, $\bm{\upsilon}(\hat{\bm{s}}_{t+1}|\bm{s}_{t},\bm{u}_{t},\bm{\theta_{\upsilon}})$, memory buffer, $\mathcal{L}_{\bm{\pi}}$, $\alpha_{\bm{\pi}}$, warmup steps $T_{\text{warm}}$, unroll steps $T_{\text{unroll}}$
@@ Appendix I Adaptive Leaky Integrate-and-Fire Neuron Model
In the following subsections, we examine how each mechanism influences the behavior of single ALIF neurons and their firing responses to tonic and noisy inputs, as well as their impact on learning the 2D control task. To allow both $\vartheta_{a}$ and $\vartheta_{b}$ to contribute meaningfully, in the control experiments we use an unroll window $T_{\text{unroll}}=40$ environment steps (= 280 SNN steps) for both the prediction and the policy model.
@@ I.3 Combined Behavior
Overall, low-rank factorization provides a practical method for scaling SNN models while controlling memory and compute requirements. This approach also aligns with biological findings suggesting that population activity often resides on low-dimensional manifolds (Averbeck et al., 2006; Churchland et al., 2012).
@@
Appendix O SNN Ablation Study and Component Analysis
To isolate which architectural components are responsible for the performance and stability gains of the proposed Pred-control SNN, we conduct a structured ablation study on the 3D reaching task. All variants are trained and evaluated under the same protocol as the full model described in the main text; only the network architecture is modified.
Ablated components and model variants.
Starting from a minimal floor model, we selectively remove or reintroduce three core components: (i) latent-state compression (Appendix N), (ii) adaptive neuron dynamics via ALIF neurons (Appendix I), and (iii) learnable membrane and synaptic time constants (Appendix H). Intermediate variants disable exactly one component at a time, enabling attribution of performance changes to individual design choices. Where applicable, we consider both a small (S, 512 neurons per layer) and a large (L, 2048 neurons per layer) configuration. All other settings—including data collection, unroll length, optimization, and loss functions—are held fixed.
Quantitative comparison.
[…]
Learning dynamics.
Learning curves in Figure 27 further reveal how individual components affect optimization. The full model converges faster, achieves higher sustained time-on-target, and exhibits smaller, more stable gradient norms. Removing ALIF neurons or learnable time constants leads to noisier optimization signals and reduced asymptotic control performance, whereas compression primarily influences performance through representational capacity rather than training stability.
Summary.
Together, these results demonstrate that the gains of the proposed SNN controller do not arise from parameter count alone. Adaptive neuron dynamics—via ALIF neurons and learnable time constants—play a central role in stabilizing training and enabling effective temporal representations, while compression provides a favorable trade-off between model size and performance. The full Pred-control SNN therefore occupies a markedly better point on the performance–efficiency frontier than any ablated variant.
Table 3: Ablation of key SNN components on the 3D reaching task. Models are evaluated at the final training iteration (means over 10 seeds; std omitted for clarity). (L) denotes the large variant with 2048 neurons per layer, and (S) denotes the small variant with 512 neurons per layer. The Basic SNN disables compression, learnable time constants, and ALIF neurons.
[…]
Neurons / layer 2048 512 2048 512 2048 2048 512
Policy parameters 315,422 78,878 4,247,582 275,486 311,326 303,122 272,402
Predictor parameters 591,914 148,010 8,456,234 541,226 587,818 579,608 538,136
Total parameters 907,336 226,888 12,703,816 816,712 899,144 882,730 810,538
Cumulative distance $\downarrow$ 0.634 0.833 0.607 0.769 0.712 0.714 0.884
Success rate $\uparrow$ 0.986 0.964 0.989 0.986 0.992 0.954 0.975
Time on target $\uparrow$ 0.615 0.364 0.652 0.435 0.492 0.559 0.322
State MSE $\downarrow$ 0.020 0.023 0.017 0.020 0.025 0.026 0.023
Figure 27: SNN ablation study on the 3D reaching task. Learning curves for Pred-control SNN variants in which individual architectural components are removed. Shown are a minimal basic model (no learnable time constants, no ALIF neurons, no parameter compression; 512 neurons per layer), the full model (learnable time constants, ALIF neurons, and compression; 2048 neurons per layer), and intermediate variants where exactly one component is ablated. Top row: task performance metrics (cumulative distance, success rate, time on target, prediction MSE). Middle row: optimization losses and gradient norms for predictor and policy. Bottom row: population activity statistics. Curves show mean over 3 seeds (sd omitted for visual clarity). The full model consistently achieves faster convergence, higher time on target, and more stable gradients. Removing adaptive neuron dynamics (learnable time constants or ALIF neurons) degrades both learning efficiency and final control quality, while compression primarily affects performance through model capacity rather than training stability.
Appendix P Robustness to target perturbations
To probe robustness beyond the nominal evaluation protocol, we performed a simple closed-loop perturbation test in which the target position is shifted whenever the end-effector enters the success radius. As shown in Figure 28, the pretrained Pred-Control SNN rapidly reorients toward the updated target and typically re-acquires it within a few control cycles, producing repeated decreases in distance-to-target after each perturbation. This behavior suggests that the learned policy does not merely memorize a fixed trajectory, but implements a reactive feedback strategy that can adapt online to changing goal states under the learned predictive–control loop. Importantly, this adaptation occurs without any online weight updates, indicating that robustness arises from the learned internal dynamics and state-dependent control mapping.
Figure 28: Robustness to within-episode target shifts. Evaluation of a pretrained Pred-Control SNN under a simple target-perturbation protocol. Whenever the end-effector enters the success region, the target position is shifted to a new location (multiple times within a single 200-step episode). Top: Example 3D end-effector trajectories under three representative episodes, illustrating repeated re-targeting behavior. Bottom: Corresponding distance-to-target traces; each increase reflects a target shift, followed by rapid error reduction as the controller re-acquires the new target. No online learning is performed during this test.
Appendix Q Non-spiking RNN Baseline and Model Size Sweep
To contextualize the proposed spiking controller, we construct a non-spiking recurrent baseline for the 3D reaching task that matches the SNN training pipeline as closely as possible. We reuse the same data collection, episode buffer, unrolling/BPTT loop, optimizer settings, losses, and action scaling; only the network implementation differs (standard floating-point PyTorch layers instead of spiking modules).
Architecture.
The baseline retains the same two-module structure (recurrent predictor + feedforward policy) and runs at the environment control rate (one recurrent step per control cycle) with the same unroll lengths as the SNN. As in the spiking setup, the predictor outputs a state increment $\Delta\hat{\bm{s}}_{t}$ and forms the next-state estimate as $\hat{\bm{s}}_{t+1}=\bm{s}_{t}+\Delta\hat{\bm{s}}_{t}$ under teacher forcing, or $\hat{\bm{s}}_{t+1}=\hat{\bm{s}}_{t}+\Delta\hat{\bm{s}}_{t}$ during autoregressive rollouts. We use an Elman RNN with ReLU nonlinearities for the predictor and the same bounded-action transformation for the policy (including $\tanh$). No parameter compression (Appendix N) is applied.
Training objective and regularization.
The predictor is trained with the same supervised one-step objective as in the SNN pipeline, i.e., predicting $\bm{s}_{t+1}$ from $(\bm{s}_{t},\bm{u}_{t})$ under teacher forcing. The policy is trained via differentiable rollouts through the frozen predictor using the same multi-step objective; we also apply the same auxiliary penalty on the policy pre-$\tanh$ activation to reduce saturation-induced gradient collapse. In the results below, we distinguish between the one-step prediction objective (Pred. loss) and the multi-step, autoregressive rollout error (Pred. MSE), which can diverge even when one-step prediction is accurate.
Model size sweep.
We vary the number of neurons per layer $\in\{32,128,512,2048\}$ and set all hidden layers in both predictor and policy to this width, keeping all other hyperparameters fixed. Table 4 summarizes the resulting parameter counts and final 3D control performance (means over 3 seeds), while Figure 29 reports learning curves over training.
Overall, control performance improves strongly with model size: larger baselines converge faster, reach higher success rates, spend more time on target, and reduce cumulative distance more quickly. The one-step prediction objective remains well-behaved across sizes (low Pred. loss), but the largest model exhibits an outlier in long-horizon autoregressive prediction accuracy: while it achieves the best control metrics, its Pred. MSE spikes late in training for an outlier run (see Figure 29), which is surprising for the large model under constant teacher forcing. Since the same training protocol (including constant teacher forcing during predictor training) is used for both baseline and SNN, we report this behaviour as-is. Policy/task performance remains unaffected by this spike.
Table 4: Non-spiking RNN baseline on the 3D reaching task. Models are evaluated at the final training iteration (means over 3 seeds; std omitted for clarity). All hidden layers in both predictor and policy use the stated width (neurons / layer). Note that Pred. MSE is a multi-step, autoregressive rollout error and can show outlier behaviour even when one-step prediction remains accurate (cf. Figure 29).
Non-spiking baseline model
Neurons / layer 32 128 512 2048
Policy parameters 1,868 19,724 275,468 4,247,564
Predictor parameters 3,219 37,395 542,739 8,462,355
Total parameters 5,087 57,119 818,207 12,709,919
Cumulative distance $\downarrow$ 2.479 1.129 0.794 0.691
Success rate $\uparrow$ 0.377 0.864 0.969 0.970
Time on target $\uparrow$ 0.031 0.226 0.492 0.601
Pred. MSE $\downarrow$ 0.144 0.074 0.013 18.587
Figure 29: Non-spiking RNN baseline (3D task): effect of model size. Learning curves for neurons per layer $\in\{32,128,512,2048\}$ applied to all hidden layers in both predictor and policy (3 seeds; mean $\pm$ s.e.m.). Control performance improves with model size across runs (higher success and time on target, lower cumulative distance), with the 512–2048 models converging fastest. While the one-step prediction objective (Pred. loss) remains low across sizes, the largest model shows an outlier in long-horizon autoregressive prediction error (Pred. MSE), illustrating that multi-step rollouts can degrade despite strong one-step fits.
Note on policy recurrence.
We also tested a recurrent policy variant under the same setup, but it did not improve performance and often reduced training stability; we therefore proceed with the feedforward policy baseline.
Appendix R Hyperparameters
@@
Unroll steps (prediction/policy) $T_{\text{unroll}}^{\upsilon},T_{\text{unroll}}^{\pi}$ 40 / 40
@@
Synaptic time constant $\tau_{\text{syn}}$ 5 ms (learned)
Adaptation time constant $\tau_{\text{ada}}$ 100 ms (learned)